What is the price of a European put option on a non-dividend paying stock when the stock price is $69, the strike price is $70, the risk-free interest rate is 5% per annum, the volatility is 35%per annum, and the time to maturity is six months? Please give me step by step by step instructions.
Answer #1

In this case, S = 69, K=70, r= 0.05 and σ= 0.35 and T=0.5

D1 = (Log (69/70)+(0.05+0.35*0.35)x0.5)/0.35x0.5√0.5 = 0.1666

D2 = d1-0.35√0.5 =-0.0809

Price of European put option

70e-0.05×0.5N(0.0809)-69N(-0.1666)

                                                =70e-0.025×0.5323-69×0.4338

=$6.4